Sun Solar has earnings before interest and taxes of $106 million and considering making a change to its capital structure toCCost of Debt = 2%
Answer #1

1). Unlevered cost of equity (RsU) = Rf + beta*Rmp where Rf (risk-free rate) = 6%; Rmp (market risk premium) = 5%

Unlevered beta = (RsU - Rf)/Rmp = (12%-6%)/5% = 1.20

Levered beta = unlevered beta*[1 + (1-Tax rate)*D/E] = 1.20*[1 + (1-40%)*0.6] = 1.63

Levered cost of equity (RsL) = Rf + beta*Rmp = 6% + (1.63*5%) = 14.16%

Cost of capital = (We*RsL) + (Wd*Rd*(1-Tax rate)) where

We (weight of equity) = 1/(1+D/E) = 1/(1+0.6) = 0.625; Wd (weight of debt) = 1- Wd = 1-0.625 = 0.375; Rd (cost of debt) = 2%

Cost of capital = (0.625*14.16%) + (0.375*2%*(1-40%)) = 9.30%

b).

Formula Rd*(1-T) bu*[1 + (1-T)*Wd/We] Rf + (B*Rmp) (Wd*After-tax cost of debt)+(We*RsL)
Percent financed with debt (Wd) Percent financed with equity (We) Bond rating Before-tax cost of debt (Rd) After-tax cost of debt Unlevered cost of quity (RsU) Unlevered beta (bu) Levered beta (B) Levered cost of equity (RsL) WACC
0.1 0.9 AAA 7.00% 4.20% 12.00% 1.20 1.28 12.40% 11.58%
0.2 0.8 AA 7.20% 4.32% 12.00% 1.20 1.38 12.90% 11.18%
0.3 0.7 A 8.00% 4.80% 12.00% 1.20 1.51 13.54% 10.92%
0.4 0.6 BBB 8.80% 5.28% 12.00% 1.20 1.68 14.40% 10.75%
0.5 0.5 BB 9.60% 5.76% 12.00% 1.20 1.92 15.60% 10.68%
0.6 0.4 B 11.00% 6.60% 12.00% 1.20 2.28 17.40% 10.92%